OpenStudy (anonymous):

solve 1 over 10 / 2 over 5 =? show your work.

OpenStudy (anonymous):

youre cross multiplying again for this one..

OpenStudy (anonymous):

is the question like this\[\frac{ 1 }{ 10 }\div \frac{ 2 }{ 5 }\]

OpenStudy (anonymous):

yes it is.

OpenStudy (anonymous):

ok then because you are divividing fractions you will need to change the divide to multiply adn flip the second fraction around

OpenStudy (anonymous):

\[\frac{ 1 }{ 10 }\div \frac{ 2 }{5 }\] is the same as saying \[\frac{ 1 }{10 }\times \frac{ 5 }{ 2 }\] Then you just straight multiply.

OpenStudy (anonymous):

so it will be \[\frac{ 1 }{ 10 }\times \frac{ 5 }{ 2 }\]

OpenStudy (anonymous):